OpenStudy (stormswan):

Please help. I will give a medal~ Prove that a line that divides two sides of a triangle proportionally is parallel to the third side. Be sure to create and name the appropriate geometric figures.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Given: \(\triangle ABC\) \(~~~~~~~~~~\dfrac{AD}{DB} = \dfrac{AE}{EC}\) Prove: \(\overline{DE} \parallel \overline{BC}\)

OpenStudy (mathstudent55):

This is the proof you need to do.

OpenStudy (stormswan):

Okay, but I don't know how to do it. Can you please show me how?

OpenStudy (mathstudent55):

You need a 2-column proof?

OpenStudy (stormswan):

ummm.. i don't think so? i think i just have to prove it.

OpenStudy (mathstudent55):

Ok, then I will give you the outline of the proof. We are given \(\dfrac{AD}{DB} = \dfrac{AE}{EC} \) Through a property of proportions, we can say \(\dfrac{AD}{AD + DB} = \dfrac{AE}{AE + EC} \) By the segment addition postulate, AD + DB = AB, and AE + EC = AC

OpenStudy (mathstudent55):

If you are following what I am writing, notice that I just corrected what I wrote just above.

OpenStudy (mathstudent55):

By substitution, \(\dfrac{AD}{AB} = \dfrac{AE}{AC} \) The proportion above shows that the lengths of the sides of triangles ADE and ABC are proportional. Also, we know that angle A is congruent to angle A. Then by the SAS Similarity theorem, triangles ABC and ADE are similar.

OpenStudy (stormswan):

Okay, that makes sense.

OpenStudy (mathstudent55):

Since those triangles are similar, then by the definition of similar triangles, angles ADE and ABC are congruent. Finally, lines DE and BC are parallel because of the postulate: "If two lines are cut by a transversal such that corresponding angles are congruent, then the lines are parallel."

OpenStudy (mathstudent55):

sorry, but gtg

OpenStudy (stormswan):

thank you! this helped a lot! :)

OpenStudy (mathstudent55):

you're welcome